OpenStudy (anonymous):

Couldn't find a root for P(x)=2x^4-x^3-14x+17...asking for all rational roots for P(x)=0

OpenStudy (mertsj):

Did you try synthetic division with all the possible rational roots?

OpenStudy (anonymous):

Yes

OpenStudy (mertsj):

Did you remember to put in the 0 for 0x^2 ??

OpenStudy (anonymous):

Lol! ty

OpenStudy (mertsj):

However, it has no rational roots: http://www.wolframalpha.com/input/?i=Solve%3A2x^4-x^3-14x%2B17%3D0

OpenStudy (anonymous):

I also noticed that i wrote 17 instead of 7.

OpenStudy (mertsj):

Well, that's a whole different thing.

OpenStudy (mertsj):

Now you can factor by grouping.

OpenStudy (anonymous):

How do you do that?

OpenStudy (campbell_st):

I think if you graph the curve you'll find its positive definite

OpenStudy (campbell_st):

so it won't have any rational roots

OpenStudy (mertsj):

\[2x^4-x^3-14x+7=x^3(2x-1)-7(2x-1)=(2x-1)(x^3-7)=0\]

OpenStudy (anonymous):

We didn't learn the graphing part for this lesson yet

OpenStudy (campbell_st):

is it +17 or +7..?

OpenStudy (mertsj):

So: \[2x-1=0; x=\frac{1}{2}\]

OpenStudy (mertsj):

@campbell_st If you read the previous posts, your question will be answered

OpenStudy (mertsj):

And that is the only rational root.

OpenStudy (mertsj):

Did you understand the "factor by grouping" part?

OpenStudy (anonymous):

Yes, you took out from the paranthesis so it became x^3(2x.... etc

OpenStudy (mertsj):

very good.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Quick question, when using synthetic division, lets say you're using a 2 to multiply into the following numbers, do you turn that 2 into a negative 2?

OpenStudy (mertsj):

If 2 is a 0, 2 is the shelf number If x+2 is a factor, -2 is the shelf number.

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

When using synthetic division does the 0 always have to be at the end? or can it be in the middle ?

OpenStudy (mertsj):

It has to be at the end, because that means that the remainder of the division problem is 0. Or, in other words, you have found a factor because the problem "came out even"

OpenStudy (anonymous):

kk ty

OpenStudy (mertsj):

yw